Q-By what percentage should the pressure of a given mass of gas be increased,so as to decrease its volume  by 10% at a constant temperature.

Grade:11

2 Answers

Akash Kumar Dutta
11 years ago

Dear Nimish,
let the % be x,
Then by applying Boyles Law, we get,
P1V1=P2V2
=> P.V=(P+x.P).(V-V/10)
=> P.V=P(1+x).V.9/10
=> 10/9=1+x
=> x=1/9 % (ANS)
Regards.

Rohan Das
11 years ago

here

P1V1/T1=P2V2/T2

BUT TEMP. REMAINS CONSTANT.

SO P1V1=P2V2

OR P1V1=P2(V1-V1/10)

OR P1V1=P2*9V1/10

OR P1=9P2/10

OR P2=10P1/9

NOW, PERCENTAGE INCREASE=(P2-P1)*100/P1

WHICH IS EQUAL TO 100/9%